Please Help
In a recent survey 64% of the cars entering the Maine turnpike at Kittery had out of state plates. Of those with out of state plates, 75% said that they were tourists. If one car is chosen at random, what is the probability that:
a)The car has Maine plates?
b)It has out of state plates but still does not contain tourists?
c)It has out of state plates and tourists?
thank you

Found 2 solutions by edjones, ewatrrr:
Answer by edjones(7569)

a) 1-.64=.36
.
b) .64*.25=.16
.
c) .64*.75=.48
.
Ed

Answer by ewatrrr(10682)


Hi,
In a recent survey 64% of the cars entering the Maine turnpike at Kittery had out of state plates, of which 75% were tourists.
P(out of state) = 64/100 = 16/25
a)The car has Maine plates?
1- 16/25 = 9/25
b)It has out of state plates but still does not contain tourists?
16% or 16/100 = 4/25
c)It has out of state plates and tourists?
48% or 48/100 = 12/25